OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −(x + 6) y − 4 = 2(x + 6) y + 6 = −(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

i think its the second one but i am not sure.

OpenStudy (anonymous):

isn't the slope -1/2?

OpenStudy (anonymous):

yea for the first and 3rd answers there is a -1/2 half for slope idk y it got deleted

OpenStudy (anonymous):

I would say that it should be y-4=-1/2(x+6)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob :)

OpenStudy (shamim):

the formula is\[y-y _{2}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-x _{2})\]

OpenStudy (shamim):

here \[x _{1}=-6 ,y _{1}=4,x _{2}=2,y _{2}=0\]

OpenStudy (shamim):

or, we can let\[x _{1}=2,y _{1}=0,x _{2}=-6,y _{2}=4\]

OpenStudy (shamim):

now u can try to solve it

OpenStudy (anonymous):

thanks